A 5.00 kg bowling ball is held 70.0cm above the ground. Determine its potential energy:
a) 34.3 Joules
b) 123 Joules
c) 3.43 x 103
Joules
d) 1.23 x 104
Joules
1
Expert's answer
2020-05-25T11:18:39-0400

We can find the potential energy of the bowling ball from the formula:


"PE = mgh = 5.0kg \cdot 9.8 \dfrac{m}{s^2} \cdot 0.7m = 34.3 J."

Answer:

"a) 34.3 J."